OpenStudy (anonymous):

x=0 is a regular singular point of the given DE. Show that the indicial roots of the singularity do not differ by an integer. Use the the method of Frobenius to obtain two linearly independent series solutions about x=0. Form the general solution on (0,\infty)

OpenStudy (anonymous):

\[2xy''-y'+2y=0\]

OpenStudy (anonymous):

frobenius theorem states \[(x-x_0)^r\sum_{n=0}^\infty c_n(x-x_0)^n=\sum_{n=0}^\infty c_n(x-x_0)^{n+r}\]

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

Question when they take the derivatives the place at which the sums start are all n=0 however when you did it before they went up one as you took the derivative

OpenStudy (anonymous):

is all n=0 since it's arbitrary until we know what r =

OpenStudy (anonymous):

@zepp

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

in other words if you had \[y=\sum_{n=0}^\infty c_nx^{n-1}\] if you plug it in you get \[x^{-1}\] which would not be a power series correct?

OpenStudy (anonymous):

so they move it up one whereas here we have n+r-1 but we do not know what r is

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

\[y=\sum_{n=0}^\infty c_nx^{n+r}\] \[y'=\sum_{n=0}^\infty c_n(n+r)x^{n+r-1}\]

OpenStudy (anonymous):

\[y''=\sum_{n=0}^\infty c_n(n+r-1)(n+r)x^{n+r-2}\]

OpenStudy (anonymous):

\[2x\sum_{n=0}^\infty c_n (n+r-1)(n+r)x^{n+r-2}-\sum_{n=0}^\infty c_n(n+r)x^{n+r-1} \] \[+2\sum_{n=0}^\infty c_nx^{n+r}\]

OpenStudy (anonymous):

\[\sum_{n=0}^\infty2c_n(n+r-1)(n+r)x^{n+r-1}-\sum_{n=0}^\infty c_n(n+r)x^{n+r-1}\] \[+\sum_{n=0}^\infty 2c_nx^{n+r}\]

OpenStudy (anonymous):

\[\sum_{n=0}^\infty c_n(2n+2r-2)(n+r)x^{n+r-1}-\sum_{n=0}^\infty c_n(n+1)x^{n+r-1}\]

OpenStudy (anonymous):

\[\sum_{n=0}^\infty [(2n+2r-2)(n+r)-n+r]c_nx^{n+r-1}\]

OpenStudy (anonymous):

for last sum let \[k=n+1\] \[k-1=n\] and first added sums n=k \[\sum_{k=0}^\infty [2k+2r-2)(k+r)-k+r]c_k x^{k+r-1}+\sum_{k=1}^\infty2c_{k-1}x^{k+r-1}\]

OpenStudy (experimentx):

is not so fond of power series solution. I can provide answer from Maple if you want.

OpenStudy (anonymous):

pull out nonzero termof first sum \[[(2r-2)(r)+r]c_0x^{r-1}+\sum_{n=1}^\infty(2k+2r-2)(k+r)-k+r]c_kx^{k+r-1}\] \[+\sum_{k=1}^\infty 2c_{k-1}x^{k+r-1}\]

OpenStudy (anonymous):

gah i need to resub -.-

OpenStudy (anonymous):

these problems are driving me a-wall